# Randomly punctured Reed–Solomon codes achieve list-decoding capacity over linear-sized fields

### v5

@@ 1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$ for $\ell=1,\dots,j$ do
3             $\tau\leftarrow$ the type of $i_{\ell-1}$ $s\leftarrow$ number of indices among $i_{\ell_{0}},i_{\ell_{0}+1},\dots,i_{\ell-1}$ that are type $\tau$ $i_{\ell-1}^{\prime}\leftarrow$ the $s$-th smallest element of $B$ that has type $\tau$ if $i_{\ell-1}^{\prime}$ is defined then
@@
7             $B\leftarrow\emptyset$ for $\tau=1,\dots,2^{t}$ do
@@
            $M_{\ell}\leftarrow$ lexicographically smallest nonsingular $(t-1)k\times(t-1)k$ submatrix of $\mathsf{RIM}_{\mathcal{H}}^{B\cup\{i_{1},\dots,i_{\ell-1}\}}$ $\ell_{0}\leftarrow\ell$ // new refresh index
9
10
@@
2       $M_{1},\dots,M_{j}=\mathtt{GetMatrixSequence}(i_{1},\dots,i_{j-1})$ $i_{j}\leftarrow$ smallest index $i$ such that $M_{j}(X_{\leq i}=\alpha_{\leq i})$ is singular if $i_{j}$ not defined then
@@
3.3 $\mathtt{GetCertificate}$ and $\mathtt{GetMatrixSequence}$: Basic properties
@@
Furthermore, $\mathtt{GetMatrixSequence}$ is a deterministic function of $i_{1},\dots,i_{j-1}$, and it computes $M_{\ell}$ “online”, meaning $M_{\ell}$ depends only on $i_{1},\dots,i_{\ell-1}$ for all $\ell=1,\dots,j$ (and $M_{1}$ is always the same matrix). In particular, $\mathtt{GetMatrixSequence}(i_{1},\dots,i_{j-1})$ is a prefix of $\mathtt{GetMatrixSequence}(i_{1},\dots,i_{j})$.
@@ 3.4 Bad evaluation points admit certificates
Here, we establish Lemma 3.8, that if some evaluation points make $\mathsf{RIM}_{\mathcal{H}}$ not full column rank, then $\mathtt{GetCertificate}$ outputs a certificate. To do so, we first justify our matrix constructions, showing that the matrices in $\mathtt{GetMatrixSequence}$ are in fact submatrices of $\mathsf{RIM}_{\mathcal{H}}$.
@@ 3.5 Bounding the number of possible certificates
Suppose for contradiction that $i_{\ell}<i_{\ell-1}$. (Note that $i_{\ell-1}\neq i_{\ell}$ by Lemma 3.6.) We contradict the first item by showing, using the second item, that $M_{\ell-1}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ is also singular. By the definition of $\mathtt{GetMatrixSequence}$, since $\ell$ is not a refresh index, $M_{\ell}$ is defined in Line 1. By construction of $B$ and $i_{\ell-1}^{\prime}$, we know that $i_{\ell-1}^{\prime}>i_{\ell-1}>i_{\ell}$. Thus, not only is $M_{\ell}$ obtained from $M_{\ell-1}$ by replacing all copies of $X_{i_{\ell-1}}$ with $X_{i_{\ell-1}^{\prime}}$, but $M_{\ell}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ is also obtained by replacing all copies of $X_{i_{\ell-1}}$ with $X_{i_{\ell-1}^{\prime}}$ in $M_{\ell-1}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ Moreover, the variable $X_{i_{\ell-1}^{\prime}}$ does not appear in $M_{\ell-1}$ by Lemma 3.5. So we conclude that, as $M_{\ell}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ is singular, so is $M_{\ell-1}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$.
@@
The number of possible outputs to $\mathtt{GetCertificate}$ is at most $\binom{n}{r}2^{tr}$.
@@ 3.6 Bounding the probability of one certificate
The goal of this section is to establish Corollary 3.12, which states that the probability of obtaining a particular certificate is at most $(\frac{(t-1)k}{q-n})^{r}$. The argument is implicit in [GZ23], but we include a proof for completeness.
Lemma 3.11 (Implicit in [GZ23]).
@@
The probability that all the events hold is at most $(\frac{(t-1)k}{q-n})^{r}$.
@@
Note that the set of evaluation points $\alpha_{1},\dots,\alpha_{n}$ for which events $E_{j}$ and $F_{j}$ occur depends only on $M_{j}$ and $i_{j}$. Furthermore, each of the events $E_{j}$ and $F_{j}$ depends only on $M_{i}$, $i_{j}$, and the evaluation points. Thus, by relabeling the index $j$, we may assume without loss of generality that $i_{1}<i_{2}<\cdots<i_{r}$. We emphasize that we are not assuming that the output of $\mathtt{GetCertificate}$ satisfies $i_{1}<\cdots<i_{r}$ (this is not true). We are instead just choosing how we ’reveal’ our events $E_{j}$ and $F_{j}$: starting with the smallest index in $i_{1},\ldots,i_{r}$ and ending with the largest index in it.
@@
Note that $E_{1}\wedge F_{1}\wedge\cdots\wedge E_{j-1}\wedge F_{j-1}\wedge E_{j}$ depends only on $\alpha_{1},\dots,\alpha_{i_{j}-1}$, and $F_{j}$ depends only on $\alpha_{1},\dots,\alpha_{i_{j}}$. For any $\alpha_{1},\dots,\alpha_{i_{j}-1}$ for which $E_{1}\wedge F_{1}\wedge\cdots\wedge E_{j-1}\wedge F_{j-1}\wedge E_{j}$ holds, we have that $M_{j}(X_{\leq i_{j}-1}=\alpha_{\leq i_{j}-1})$ is a $(t-1)k\times(t-1)k$ matrix in $\mathbb{F}_{q}(X_{i_{j}},X_{i_{j}+1},\dots,X_{n})$ whose determinant is a nonzero polynomial of degree at most $(t-1)k$ in each variable (the determinant contains at most $t-1$ rows including $X_{i_{j}}$, each time with maximum degree $k-1$). In particular, at most $(t-1)k$ values of $\alpha_{i_{j}}$ can make the determinant zero since, viewing the determinant as a polynomial in variables $X_{i_{j}+1},\dots,X_{n}$ with coefficients in $\mathbb{F}_{q}[X_{i_{j}}]$, any single nonzero coefficient becomes zero on at most $(t-1)k$ values of $\alpha_{i_{j}}$. Conditioned on $\alpha_{1},\dots,\alpha_{i_{j}-1}$, the field element $\alpha_{i_{j}}$ is uniformly random over $q-i_{j}+1\geq q-n$ elements. Thus, we have, for all $\alpha_{1},\dots,\alpha_{i_{j}-1}$ such that $E_{1}\wedge F_{1}\wedge\cdots\wedge E_{j-1}\wedge F_{j-1}\wedge E_{j}$,
@@
Recall (Section 3.2) that we fixed $\mathcal{H}$ to be a type-ordered $(k+\varepsilon n)$-weakly-partition-connected hypergraph. By Lemma 3.8, if the matrix $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ does not have full column rank, then $\mathtt{GetCertificate}(\alpha_{1},\dots,\alpha_{n})$ is some certificate $(i_{1},\dots,i_{r})$. The probability that $\mathtt{GetCertificate}(\alpha_{1},\dots,\alpha_{n})=(i_{1},\dots,i_{r})$ is at most $(\frac{(t-1)k}{q-n})^{r}$ by Corollary 3.12. By Corollary 3.10, there are at most $\binom{n}{r}2^{tr}$ certificates. Taking a union bound over possible certificates gives the lemma. ∎
@@ 4.1 Preliminaries: Notation and Definitions
$\mathcal{G}$ =$\stackrel{{\scriptstyle\rm def}}{{=}}\begin{bmatrix}X_{1,1}&\cdots&X_{1,k}\\ \vdots&\ddots&\vdots\\ X_{n,1}&\cdots&X_{n,k}\\ \end{bmatrix}\in\mathbb{F}_{q}(X_{1,1},\dots,X_{n,k})^{n\times k},$ (25)
$G$ =$\stackrel{{\scriptstyle\rm def}}{{=}}\mathcal{G}|_{X_{[n]\times[n]}=\alpha_{[n]\times[k]}}$
$\mathcal{G}_{i}$ =$\stackrel{{\scriptstyle\rm def}}{{=}}[X_{i,1},\dots,X_{i,k}]\text{ (the $i$th row of $\mathcal{G}$).}$. (26)
@@
The reduced intersection matrix $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}$ associated with a hypergraph $\mathcal{H}=([t],(e_{1},\dots,e_{n}))$ is a $\operatorname*{wt}(\mathcal{E})\times(t-1)k$ matrix over the field of fractions $\mathbb{F}_{q}(X_{1,1},\dots,X_{n,k})$. For each hyperedge $e_{i}$ with vertices $j_{1}<j_{2}<\dots<j_{|e_{i}|}$, we add $\operatorname*{wt}(e_{i})=|e_{i}|-1$ rows to $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}$. For $u=2,\dots,|e_{i}|$, we add a row $r_{i,u}=(r^{(1)},\ldots,r^{(t-1)})$ of length $(t-1)k$ defined as follows:
@@ 4.2 Preliminaries: Properties of RLC Reduced Intersection Matrices
We remark that Lemma 4.3 can be proven directly by following the proof framework of Theorem 2.11 in Appendix A.3, but instead substitute the use of Theorem A.2 with an analogous GM-MDS theorem for Random Linear Codes, which can be found in Lemma 7 of [DSY15] (Lemma 7 of [DSY15] only implies Lemma 4.3 for $q$ to be sufficiently large, but again by Remark 2.12 the $q$ sufficiently large version of Lemma 4.3 implies the lemma for all $q$). That way, the proof of Theorem 1.3 relies only on the proof framework of Theorem 1.1 and not on any of its lemmas.
@@
Let $\mathcal{H}=([t],\mathcal{E})$ be a $(k+\varepsilon n)$-weakly-partition-connected hypergraph with $t\geq 2$. For all sets $B\subset[n]$ with $|B|\leq\varepsilon n$, we have that $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}^{B}$ is nonempty and has full column rank.
@@ 4.3 The proof
1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$ for $\ell=1,\dots,j$ do
@@
3             $\tau\leftarrow$ the type of $i_{\ell-1}$ $s\leftarrow$ number of indices among $i_{\ell_{0}},i_{\ell_{0}+1},\dots,i_{\ell-1}$ that are type $\tau$ $i_{\ell-1}^{\prime}\leftarrow$ the $s$-th smallest element of $B$ that has type $\tau$ if $i_{\ell-1}^{\prime}$ is defined then

### v6

@@ 1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$
4    $\tau\leftarrow$ the type of $i_{\ell-1}$
5    $s\leftarrow$ number of indices among $i_{\ell_{0}},i_{\ell_{0}+1},\dots,i_{\ell-1}$ that are type $\tau$
6    $i_{\ell-1}^{\prime}\leftarrow$ the $s$-th smallest element of $B$ that has type $\tau$
7    if $i_{\ell-1}^{\prime}$ is defined then
@@ 6    $i_{\ell-1}^{\prime}\leftarrow$ the $s$-th smallest element of $B$ that has type $\tau$
11    $B\leftarrow\emptyset$
12    for $\tau=1,\dots,2^{t}$ do
@@ 11    $B\leftarrow\emptyset$
14
15    $M_{\ell}\leftarrow$ lexicographically smallest nonsingular $(t-1)k\times(t-1)k$ submatrix of $\mathsf{RIM}_{\mathcal{H}}^{B\cup\{i_{1},\dots,i_{\ell-1}\}}$
    $\ell_{0}\leftarrow\ell$ // new refresh index
16
17
18
@@
2 $M_{1},\dots,M_{j}=\mathtt{GetMatrixSequence}(i_{1},\dots,i_{j-1})$
3 $i_{j}\leftarrow$ smallest index $i$ such that $M_{j}(X_{\leq i}=\alpha_{\leq i})$ is singular
4 if $i_{j}$ not defined then
@@ 3 $i_{j}\leftarrow$ smallest index $i$ such that $M_{j}(X_{\leq i}=\alpha_{\leq i})$ is singular
[…]
@@
$\mathtt{GetMatrixSequence}$ is a deterministic function of $i_{1},\dots,i_{j-1}$, and it computes $M_{\ell}$ “online”, meaning $M_{\ell}$ depends only on $i_{1},\dots,i_{\ell-1}$ for all $\ell=1,\dots,j$ (and $M_{1}$ is always the same matrix). In particular, $\mathtt{GetMatrixSequence}(i_{1},\dots,i_{j-1})$ is a prefix of $\mathtt{GetMatrixSequence}(i_{1},\dots,i_{j})$.
@@ 3.4 Bad Evaluation Points Admit Certificates
Here, we establish Lemma 3.8, that if some evaluation points make $\mathsf{RIM}_{\mathcal{H}}$ not full-column-rank, then $\mathtt{GetCertificate}$ outputs a certificate. To do so, we first justify our matrix constructions, showing that the matrices in $\mathtt{GetMatrixSequence}$ are in fact submatrices of $\mathsf{RIM}_{\mathcal{H}}$.
@@ 3.5 Bounding the Number of Possible Certificates
Suppose for contradiction that $i_{\ell}<i_{\ell-1}$. (Note that $i_{\ell-1}\neq i_{\ell}$ by Lemma 3.6.) We contradict the first item by showing, using the second item, that $M_{\ell-1}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ is also singular. By the definition of $\mathtt{GetMatrixSequence}$, since $\ell$ is not a refresh index, $M_{\ell}$ is defined in Line 1. By construction of $B$ and $i_{\ell-1}^{\prime}$, we know that $i_{\ell-1}^{\prime}>i_{\ell-1}>i_{\ell}$. Thus, not only is $M_{\ell}$ obtained from $M_{\ell-1}$ by replacing all copies of $X_{i_{\ell-1}}$ with $X_{i_{\ell-1}^{\prime}}$, but $M_{\ell}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ is also obtained by replacing all copies of $X_{i_{\ell-1}}$ with $X_{i_{\ell-1}^{\prime}}$ in $M_{\ell-1}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ . Moreover, the variable $X_{i_{\ell-1}^{\prime}}$ does not appear in $M_{\ell-1}$ by Lemma 3.5. So we conclude that, as $M_{\ell}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$ is singular, so is $M_{\ell-1}(X_{\leq i_{\ell}}=\alpha_{\leq i_{\ell}})$.
@@
The number of possible outputs of $\mathtt{GetCertificate}$ is at most $\binom{n}{r}2^{tr}$.
@@ 3.6 Bounding the Probability of One Certificate
The goal of this section is to establish Corollary 3.12, which states that the probability of obtaining a particular certificate is at most $\left(\frac{(t-1)k}{q-n}\right)^{r}$.
Lemma 3.11.
@@
The probability that all the events hold is at most $\left(\frac{(t-1)k}{q-n}\right)^{r}$.
@@
Note that the set of evaluation points $\alpha_{1},\dots,\alpha_{n}$ for which events $E_{j}$ and $F_{j}$ occur depends only on $M_{j}$ and $i_{j}$. Furthermore, each of the events $E_{j}$ and $F_{j}$ depends only on $M_{i}$, $i_{j}$, and the evaluation points. Thus, by relabeling the index $j$, we may assume without loss of generality that $i_{1}<i_{2}<\cdots<i_{r}$. We emphasize that we are not assuming that the output of $\mathtt{GetCertificate}$ satisfies $i_{1}<\cdots<i_{r}$ (this is not true). We are instead just choosing how we “reveal” our events $E_{j}$ and $F_{j}$: starting with the smallest index in $i_{1},\ldots,i_{r}$ and ending with the largest index in it.
@@
Note that $E_{1}\wedge F_{1}\wedge\cdots\wedge E_{j-1}\wedge F_{j-1}\wedge E_{j}$ depends only on $\alpha_{1},\dots,\alpha_{i_{j}-1}$, and $F_{j}$ depends only on $\alpha_{1},\dots,\alpha_{i_{j}}$. For any $\alpha_{1},\dots,\alpha_{i_{j}-1}$ for which $E_{1}\wedge F_{1}\wedge\cdots\wedge E_{j-1}\wedge F_{j-1}\wedge E_{j}$ holds, we have that $M_{j}(X_{\leq i_{j}-1}=\alpha_{\leq i_{j}-1})$ is a $(t-1)k\times(t-1)k$ matrix in $\mathbb{F}_{q}(X_{i_{j}},X_{i_{j}+1},\dots,X_{n})$ whose determinant is a nonzero polynomial of degree at most (t1)k(t-1)k in each variable (the determinant contains at most t1t-1 rows including $X_{i_{j}}$, each time with maximum degree k1k-1). In particular, at most (t1)k(t-1)k values of $\alpha_{i_{j}}$ can make the determinant zero since, viewing the determinant as a polynomial in variables $X_{i_{j}+1},\dots,X_{n}$ with coefficients in $\mathbb{F}_{q}[X_{i_{j}}]$, any single nonzero coefficient becomes zero on at most (t1)k(t-1)k values of $\alpha_{i_{j}}$. Conditioned on $\alpha_{1},\dots,\alpha_{i_{j}-1}$, the field element $\alpha_{i_{j}}$ is uniformly random over $q-i_{j}+1\geq q-n$ elements. Thus, we have, for all $\alpha_{1},\dots,\alpha_{i_{j}-1}$ such that $E_{1}\wedge F_{1}\wedge\cdots\wedge E_{j-1}\wedge F_{j-1}\wedge E_{j}$,
@@
Recall (Section 3.2) that we fixed $\mathcal{H}$ to be a type-ordered $(k+\varepsilon n)$-weakly-partition-connected hypergraph. By Lemma 3.8, if the matrix $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ does not have full column rank, then $\mathtt{GetCertificate}(\alpha_{1},\dots,\alpha_{n})$ is some certificate $(i_{1},\dots,i_{r})$. By Corollary 3.12, the probability that $\mathtt{GetCertificate}(\alpha_{1},\dots,\alpha_{n})=(i_{1},\dots,i_{r})$ holds is at most $\left(\frac{(t-1)k}{q-n}\right)^{r}$. By Corollary 3.10, there are at most $\binom{n}{r}2^{tr}$ certificates. Taking a union bound over possible certificates gives the lemma. ∎
3.8 Finishing the Proof of Theorem 1.1
Proof of Theorem 1.1.
By Lemma 2.3, if $\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})$ is not $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ average-radius list-decodable, then there exists a vector $y$ and pairwise distinct codewords $c^{(1)},\dots,c^{(t)}$ with $t\geq 2$ such that the agreement hypergraph $\mathcal{H}=([t],\mathcal{E})$ is $(R+\varepsilon)n=(k+\varepsilon n)$-weakly-partition-connected. By Lemma 2.8, the matrix $\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})$ is not full column rank. Now, the number of possible agreement hypergraphs $\mathcal{H}$ is at most $\sum_{t=2}^{L+1}2^{tn}\leq 2^{(L+2)n}$. Thus by the union bound over possible agreement hypergraphs $\mathcal{H}$ with Lemma 3.1, we have, for $r=\lfloor{\frac{\varepsilon n}{2}}\rfloor$,
$\mathop{\bf Pr\/}_{\alpha_{[n]}}\left[\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})\text{ is not $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ list-decodable}\right]$
[…]
$\leq 2^{(L+2)n}\max_{\text{$(k+\varepsilon n)$-w.p.c. }\mathcal{H}}\quad\mathop{\bf Pr\/}_{\alpha_{[n]}}\left[\mathsf{RIM}_{\mathcal{H}}(X_{[n]}=\alpha_{[n]})\text{ is not full column rank}\right]$
as desired. Here, we used that $q=n+k\cdot 2^{10L/\varepsilon}$. ∎
@@ 4.1 Preliminaries: Notation and Definitions
$\mathcal{G}$ $\stackrel{{\scriptstyle\rm def}}{{=}}\begin{bmatrix}X_{1,1}&\cdots&X_{1,k}\\ \vdots&\ddots&\vdots\\ X_{n,1}&\cdots&X_{n,k}\\ \end{bmatrix}\in\mathbb{F}_{q}(X_{1,1},\dots,X_{n,k})^{n\times k},$
$G$ $\stackrel{{\scriptstyle\rm def}}{{=}}\mathcal{G}|_{X_{[n]\times[n]}=\alpha_{[n]\times[k]}}$
$\mathcal{G}_{i}$ $\stackrel{{\scriptstyle\rm def}}{{=}}[X_{i,1},\dots,X_{i,k}]\text{ (the $i$th row of $\mathcal{G}$).}$ (16)
@@
The reduced intersection matrix $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}$ associated with a hypergraph $\mathcal{H}=([t],(e_{1},\dots,e_{n}))$ is a $\operatorname*{wt}(\mathcal{E})\times(t-1)k$ matrix over the field of fractions $\mathbb{F}_{q}(X_{1,1},\dots,X_{n,k})$. For each hyperedge $e_{i}$ with vertices $j_{1}<j_{2}<\dots<j_{|e_{i}|}$, we add $\operatorname*{wt}(e_{i})=|e_{i}|-1$ rows to $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}$. For $u=2,\dots,|e_{i}|$, we add a row $r_{i,u}=(r^{(1)},\ldots,r^{(t-1)})$ of length (t1)k(t-1)k defined as follows:
@@ 4.2 Preliminaries: Properties of RLC Reduced Intersection Matrices
We remark that Lemma 4.3 can be proven directly by following the proof framework of Theorem 2.11 in Appendix A.3, but instead substitute the use of Theorem A.2 with an analogous GM-MDS theorem for Random Linear Codes, which can be found in Lemma 7 of [14] (Lemma 7 of [14] only implies Lemma 4.3 for $q$ to be sufficiently large, but again by Remark 2.12 the $q$ sufficiently large version of Lemma 4.3 implies the lemma for all $q$). That way, the proof of Theorem 1.3 relies only on the proof framework of Theorem 1.1 and not on any of its lemmas.
@@
Let $\mathcal{H}=([t],\mathcal{E})$ be a $(k+\varepsilon n)$-weakly-partition-connected hypergraph with $t\geq 2$. For all sets $B\subseteq[n]$ with $|B|\leq\varepsilon n$, we have that $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}^{B}$ is nonempty and has full column rank.
@@ 4.3 The Proof
1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$
2 for $\ell=1,\dots,j$ do
@@ 1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$
4    $\tau\leftarrow$ the type of $i_{\ell-1}$
5    $s\leftarrow$ number of indices among $i_{\ell_{0}},i_{\ell_{0}+1},\dots,i_{\ell-1}$ that are type $\tau$
6    $i_{\ell-1}^{\prime}\leftarrow$ the $s$-th smallest element of $B$ that has type $\tau$
7    if $i_{\ell-1}^{\prime}$ is defined then